Graph the following equations (use points to help you graph)
y=-x+2
y=-5x-6
solution=(__,__)?

Answer by Alan3354(69443)

y=-x+2
y=-5x-6
-----------
They're straight lines so you need 2 points.
y=-x+2
Pick 2 values of x and find y, any 2 values
x = 0, y = 2 (0,2)
x = 4, y = -2 (4,-2)
Plot the 2 points, draw a line thru them.
Do the same for the other eqn
The intersection is the solution of the system.
-----------------
dl the FREE graph software at
http:www.padowan.dk
Use INS and type -x + 2, then INS and -5x - 6
---------
The solution=(__,__) is not a big butt.